35. What is the Hill Coefficient of a hemoglobin molecule that has a fractional saturation of 0.9 at 77torr? (3pts)

This is all the information
Answer #1

Log (y/1-y) = n(log pO2) - n log P50

log (0.9/1-0.9) = n ( log 26) - (log 77)

0.9542 = 0.4715 n

n = 2.02

So hill cofficient of hb molecule is 2.02.